# Hermitian and non-Hermitian topology in active matter

### v1

@@ 4.2.2 Classical analogs of quantum anomalous Hall effect
As is discussed in Sec. 3, a net (effective) magnetic field is not a requirement to realize a topological band structure. In particular, counterparts of the quantum anomalous Hall effect can also be realized in active matter. To construct such a topological active-matter system, one can utilize a kagome-lattice structure, which has been also considered in a model of the quantum Hall effect in a ferromagnetic material [274].
@@
[…]
@@
Analogs of the quantum anomalous Hall effect can also be utilized in active biological networks. For example, the game-theoretic model of competing agents on a kagome-lattice network can exhibit topological edge flows [275, 276, 277]. In more detail, one considers agents that play the rock-paper-scissors game and can move to neighbor sites. If one properly constructs a kagome-lattice network where the agents move, the agent system exhibits effective vector potentials similar to that in a quantum anomalous Hall system and thus shows a robust boundary density wave. We note that the governing equations are the coupled Lotka-Volterra equation [278], while the linear analysis around the steady state bridges such nonlinear equations, and one can employ the notion of topology as is the case in active hydrodynamics.
@@
At the end of this section, we discuss a possible application of the active-matter analogs of the quantum (anomalous) Hall effect. One prominent feature of a topological edge mode is its unidirectional propagation. Utilizing the unidirectionality, one may construct phononic counterparts of diodes (heat diodes) [130]. As is discussed above, the advantage of active metamaterials compared to passive ones is that they require no external forces to realize effective magnetic fields that lead to topologically nontrivial band structures.
@@ 4.3.1 Skin effect
In the previous section, we have focused on the Hermitian topology of active matter, while the effective Hamiltonian of active matter can naturally become non-Hermitian as can be seen from Eq. (86). Such non-Hermiticity is fairly common because active matter constantly consumes energy stored inside (e.g., animals get energy from their food and use it to move), which leads to dissipation. Utilizing such non-Hermiticity, active matter can exhibit various topological phenomena, such as skin effects and exceptional edge modes.
An experimental study shows that active matter in engineered microchannels can exhibit the second-order non-Hermitian skin effect [202]. Here, it imitates a ladder-like model of the second-order skin effect discussed above [195]. The nonreciprocal coupling is realized by using heart-shaped ratchets that are located on the way of the horizontal microchannels. Such ratchets convert the direction of motion of active particles; active particles coming from the front of the heart-shaped channel get into the channel and move along the channel’s wall. When they come out of the channel, the direction of motion will be changed. Inversely, active particles from the other side seldom get inside the channel due to the shape of the entrance. Therefore, there exist more active particles moving to the front of the heart-shaped channel than ones moving in the opposite direction, which indicates the nonreciprocal transport of active particles. The probability of moving to the neighbor lattice point depends on the width of the channels. Finally, one obtains the microchannels whose hopping amplitudes imitate those of the tight-binding model of the second-order skin effect.
The non-Hermitian skin effect has been also observed in chiral active matter without using complicated periodic structures. Cells mentioned in the previous section [268] also show the localization of the bulk modes and the modification of the dispersion relation, i.e., the non-Hermitian skin effect. One intriguing aspect of the skin effect observed here is the direction of the localization depends on the wavenumber in the other direction. This wavenumber-dependent localization can be analyzed by one directional winding number defined in each wavenumber sector. Similar wavenumber dependence is also discussed in electrical circuits [187].
@@
In practice, the non-Hermitian skin effect may play an important role in metamaterial designing and biophysics. Active metamaterials can realize nonreciprocal coupling by using active devices driven by external energy injection. In particular, a time-reversal odd response known as odd elasticity [261] has attracted interest because it can realize non-Hermitian properties in mechanical metamaterials. Localization of all the bulk modes in the skin effect implies the amplification of the response to the perturbation to one side. This amplification could be utilized in lasing devices or sensors [186]. The localization and nonreciprocal coupling are also ubiquitous in biological processes [32, 33]. We may understand such localization from a topological point of view, which might bring us a universal theory of the robustness of biological systems.
@@ 4.3.2 Exceptional point and exceptional edge modes
Exceptional points, yet another non-Hermitian topological feature, can play an important role in active-matter dynamics. One example can be found in the exceptional edge modes of chiral active matter. A study of exceptional edge modes [245] has proposed that a mixture of right- and left-rotating active particles can show exceptional edge modes. As discussed above, right- or left-rotating particles consist of counterparts of a Chern insulator for each. After introducing polar or antipolar interaction and chirality flips, the effective Hamiltonian of this mixture of chiral active matter becomes
@@
$C=\left(\begin{array}[]{ccc}i\gamma&0&0\\ 0&i\beta&0\\ 0&0&i\beta\end{array}\right).$ (91)
The alignment interaction and chirality flips lead to the non-Hermitian coupling terms like dissipative terms discussed in Sec. 3. Consequently, the topological gapless modes form a pair of exceptional points. Since chirality is broadly relevant to biological active matter, such as bacteria [62] and motility assays [90], exceptional edge modes should appear in such chiral active matter if one can realize the coexistence of left- and right-moving particles and their interactions.
Exceptional points can also affect the asymptotic behavior of the (dynamical) phase transition of active matter. If one considers chasers and preys on a circle, the eternally chasing (i.e., oscillatory) phase and the caught (i.e., static) phase can emerge depending on the parameters. A similar active system with nonreciprocal interactions can be realized in a collection of robots [96]. By analyzing such robotic active matter, the transition to the oscillatory phase accompanies exceptional points in the parameter space. The difference from the conventional phase transition emerges in, e.g., their upper critical dimension $d_{c}=8$ [287], which is different from passive systems such as the Ising model ($d_{c}=4$). Similar transitions and appearance of the exceptional points are also discussed in active Cahn-Hilliard models [288], which is originally a continuum model of the phase separation and has an additional nonreciprocal term that expresses activity.
@@ 4.4 Other activity-induced topological phenomena
Non-Hermiticity in active systems can further enrich topological phenomena associated with localized edge modes. A study on biological network systems [104] reveals that nonreciprocal hoppings reproduce a chiral edge current in a steady state of a stochastic process. We note that in a stochastic process with a Hermitian rate matrix (in other words, a purely dissipative stochastic process), a steady-state current is prohibited, while the non-Hermiticity due to the nonreciprocal hoppings induces the edge current. The emergence of the edge current can be schematically understood; once the particle arrives at the edge, the hopping to the edge site is much larger than that to the bulk site. Therefore, many particles stay around the edge and are transported along the direction of the hopping, which induces the chiral edge current.
@@ 5.1 Topological active matter in one and three dimensions
While topological active matter has been well explored in two-dimensional systems, topological phenomena should be also abundant in the other dimensions as we have discussed in Sec. 3. One possibility is a one-dimensional case, which can be realized by utilizing, e.g., molecular motors [32], microswimmers in a channel [291], and cell adhesions [143, 284]. In particular, while the skin effect is realized by using microchannels, the associated breakdown of the bulk-boundary correspondence has so far not been confirmed. Due to the nonreciprocity inherent to active matter, one-dimensional active systems should offer an ideal platform to study non-Hermitian topology.
@@ 5.2 Nonlinear topology
To derive the effective Hamiltonian of active matter, we have linearized the hydrodynamic equations of active matter for the sake of simplicity. While the effective Hamiltonian can describe the dynamics of small fluctuations of the density and velocity fields around the steady state, the whole dynamics is ultimately governed by the nonlinear hydrodynamic equations. Recently, studies on topology of nonlinear systems have been initiated given the fact that nonlinearity is fairly common in classical systems, including photonics [293] and mechanical lattices [294]. So far, studies of nonlinear topology have revealed interplay between topology and well-known nonlinear phenomena such as topological edge solitons [295] and topological synchronization [296, 297].
@@
$C=\frac{1}{2\pi i}\int_{\rm BZ}\sum_{a,b,j}\epsilon_{ab}\partial_{a}\{\left[u_{j}/||\mathbf{u}||\right]\partial_{b}\left[u_{j}/||\mathbf{u}||\right]\}dS,$ (93)
with $u_{j}$ being the $j$th component of the nonlinear eigenvector, $||\mathbf{u}||$ being the norm of the nonlinear eigenvector, $\epsilon_{ab}$ satisfying $\epsilon_{xy}=1$, $\epsilon_{yx}=-1$, and $\epsilon_{xx}=\epsilon_{yy}=0$, and $\partial_{a,b}$ being the wavenumber derivative in the $a$, $b$ direction.
The crucial difference from linear systems is that a multiple of a nonlinear eigenvector is not necessarily a nonlinear eigenvector because of the absence of the superposition principle. Said differently, a nonlinear eigenvector also depends on its norm $||\mathbf{u}||$, and one should impose $||\mathbf{u}||=w$ independently of the wavenumber $\mathbf{k}$ by introducing an additional parameter $w$ of the nonlinear eigenequation. The norm dependence of the nonlinear eigenvectors and associated nonlinear topological invariants reveals the existence of the nonlinearity-induced topological phase transition, where the existence or absence of topological edge modes is tuned by the amplitude of nonlinear waves [301]. Such amplitude dependence of topological edge modes is a genuinely nonlinear effect on topological insulators and can lead to autonomous control of topological edge modes in nonlinear systems.
@@ 5.3 Topology induced by active forces
Active elements in metamaterials can induce nonreciprocal forces that are unique to nonequilibrium systems. In particular, if the metamaterial breaks the time-reversal symmetry, one can observe an active force perpendicular to the direction of the tension, which is known as odd elasticity [261]. Another time-reversal odd response is seen in odd viscosity [8], which we have already discussed in topological chiral active matter.
Recent studies [9, 302] have proposed topological edge modes and non-Hermitian skin effects using odd elasticity. Since the breakdown of the time-reversal symmetry is ubiquitous in chiral active matter, one may utilize the active force to realize topological active matter. In fact, a recent experimental study [303] has observed signatures of odd elasticity and chiral edge current in collections of starfish embryos, while the topological origin of the chiral edge current is not clarified. Furthermore, since odd elasticity realizes symmetry-breaking and nonreciprocal interactions that are difficult to implement in conventional materials, such active forces can be useful for investigating non-Hermitian topological phenomena. It is also noteworthy that time-reversal odd responses are analogous to the Hall conductivity in condensed matter, and thus their direct interplay with nontrivial topology (e.g., quantization) should also be investigated.

### v2

@@ 5.2.2 Classical analogs of quantum anomalous Hall effect
As is discussed in Sec. 3, a net (effective) magnetic field is not a requirement to realize a topological band structure. In particular, counterparts of the quantum anomalous Hall effect can also be realized in active matter. To construct such a topological active-matter system, one can utilize a kagome-lattice structure in Fig. 21 [271], which has been also considered in a model of the quantum Hall effect [Fig. 21(b)] in a ferromagnetic material [282]. We note that the kagome lattice is a dual lattice of the honeycomb lattice, and thus the same physics explains the nontrivial topology of the kagome-lattice model and the Haldane model in Fig. 5. The theoretical proposal of an active-matter counterpart of the quantum anomalous Hall effect utilized the experimental technique to control bacterial turbulence by poles (discussed in Sec. 2.1.2), and rectify the steady-state flow so that it can imitate the vector potential in the kagome-lattice model of the quantum anomalous Hall effect. Then, it was numerically shown that the linearized dynamics (Eqs. (85) and (86)) of the proposed system exhibit gapless edge modes.
@@
$\int_{\Omega}(\nabla\times\mathbf{V}_{\rm ss})\cdot d\mathbf{S}=\oint_{\partial\Omega}\mathbf{V}_{\rm ss}\cdot d\mathbf{r},$ (90)
@@
Analogs of the quantum anomalous Hall effect can also be utilized in active biological networks. For example, the game-theoretic model of competing agents on a kagome-lattice network can exhibit topological edge flows [283, 284, 285]. In more detail, one considers agents that play the rock-paper-scissors game and can move to neighbor sites. If one properly constructs a kagome-lattice network where the agents move, the agent system exhibits effective vector potentials similar to that in a quantum anomalous Hall system and thus shows a robust boundary density wave. We note that the governing equations are the coupled Lotka-Volterra equation [286], a nonlinear equation of a particle-number distribution as the rate equation of a biochemical network (3), while the linear analysis around the steady state bridges such nonlinear equations, and one can employ the notion of topology as is the case in active hydrodynamics.
@@
At the end of this section, we discuss a possible application of the active-matter analogs of the quantum (anomalous) Hall effect. One prominent feature of a topological edge mode is its unidirectional propagation (cf. Fig. 4(a)). Utilizing the unidirectionality, one may construct phononic counterparts of diodes (heat diodes) [133]. As is discussed above, the advantage of active metamaterials compared to passive ones is that they require no external forces to realize effective magnetic fields that lead to topologically nontrivial band structures.
@@ 5.3.1 Skin effect
In the previous section, we have focused on the Hermitian topology of active matter, while the effective Hamiltonian of active matter can naturally become non-Hermitian as can be seen from Eq. (89). Such non-Hermiticity is fairly common because active matter constantly consumes energy stored inside (e.g., animals get energy from their food and use it to move), which leads to dissipation. Utilizing such non-Hermiticity, active matter can exhibit various topological phenomena, such as skin effects (cf. Sec. 4.3) and exceptional edge modes (cf. Sec. 4.4).
An experimental study shows that active matter in engineered microchannels [Fig. 22(a)] can exhibit the second-order non-Hermitian skin effect [210]. Here, it imitates a ladder-like model of the second-order skin effect discussed in Sec. 4.3.2 [203]. The nonreciprocal coupling is realized by using heart-shaped ratchets that are located on the way of the horizontal microchannels. Such ratchets convert the direction of motion of active particles; active particles coming from the front of the heart-shaped channel get into the channel and move along the channel’s wall. When they come out of the channel, the direction of motion will be changed. Inversely, active particles from the other side seldom get inside the channel due to the shape of the entrance. Therefore, there exist more active particles moving to the front of the heart-shaped channel than ones moving in the opposite direction, which indicates the nonreciprocal transport of active particles. The probability of moving to the neighbor lattice point depends on the width of the channels. Finally, one obtains the microchannels whose hopping amplitudes imitate those of the tight-binding model of the second-order skin effect. In real experiments, they observed the distribution of the active colloids (cf. Sec. 2.1.4) in the micro channels and confirmed its localization to the corner of the system shown in Fig. 22(b).
A very recent theoretical work [291] has also discussed further localization of chiral edge modes by the non-Hermitian skin effect in a nonreciprocal kagome network. They have analyzed such active fluid in a nonreciprocal network by using a transfer matrix and Floquet (i.e., time-periodic-system) techniques [292] and found that chiral edge states characterized by the Chern number or a Floquet topological invariant can have a larger amplitude at one corner than they do at the others.
The non-Hermitian skin effect has been also observed in chiral active matter without using complicated periodic structures. Cells mentioned in the previous section [275] also show the localization of the bulk modes and the modification of the dispersion relation, i.e., the non-Hermitian skin effect. One intriguing aspect of the skin effect observed here is the direction of the localization depends on the wavenumber in the other direction. This wavenumber-dependent localization can be analyzed by one directional winding number (cf. Eq. (45)) defined in each wavenumber sector. Similar wavenumber dependence is also discussed in electrical circuits [194].
@@
In practice, the non-Hermitian skin effect may play an important role in metamaterial designing and biophysics. Active metamaterials can realize nonreciprocal coupling by using active devices driven by external energy injection. In particular, a time-reversal odd response known as odd elasticity [269] has attracted interest because it can realize non-Hermitian properties in mechanical metamaterials. Localization of all the bulk modes in the skin effect implies the amplification of the response to the perturbation to one side. Thereby, this amplification would be utilized in lasing devices [297] or sensors [298]. As an example, in a nonreciprocal mechanical metamaterial [193], Ref. [297] has shown a highly large response to perturbations due to nonorthogonality of the topological edge states. Since the nonorthogonality also emerges at singularity points on the generalized Brillouin zone [299], it would be intriguing to investigate a non-Hermitian response from the viewpoint of the non-Bloch band theory (cf. Sec. 4.3.3). The localization and nonreciprocal coupling are also ubiquitous in biological processes [30, 31], which we have discussed in Sec. 2.2. We may understand such localization from a topological point of view, which might bring us a universal theory of the robustness of biological systems.
Although recent intensive studies have shed light on the point-gap topology leading to the non-Hermitian skin effect in active systems, a comprehensive understanding of the line-gap topology is still an open problem. Pioneering works have theoretically proposed that metamaterials with odd elasticity can realize topological edge states under the non-Hermitian skin effect [300, 301]. Remarkably, Ref. [300] has shown that, in a rotor chain emulating the SSH model, the winding number defined from the generalized Brillouin zone (cf. Sec. 4.3.3) can predict the existence or absence of topological edge states. The previous work has also applied a similar non-Bloch framework to a 2D rotor lattice emulating the Haldane model to predict the topological edge states. However, the relation between the framework and the amoeba formulation (cf. Sec. 4.3.4) remains unclear.
@@ 5.3.2 Exceptional point and exceptional edge modes
Exceptional points, yet another non-Hermitian topological feature, can play an important role in active-matter dynamics. One example can be found in the exceptional edge modes of chiral active matter. A study of exceptional edge modes [124] has proposed that a mixture of right- and left-rotating active particles in Fig. 23(a) can show exceptional edge modes. As discussed above, right- or left-rotating particles consist of counterparts of a Chern insulator for each.
To discuss the existence of exceptional edge modes in such a mixture of chiral active matter, Ref. [124] has first considered a particle-based model of chiral active particles. It introduces polar or antipolar interaction and chirality flips and assumes that the particles are dilute and thus three or more particle collisions do not occur. Then, the stochastic dynamics of the model is described as
$\mathbf{r}_{j}(t+dt)$ $=$ $\mathbf{r}_{j}(t)+v_{0}dt\mathbf{n}(\theta_{j}),$ (91)
$\theta_{j}(t+dt)$ $=$ $\left\{\begin{array}[]{c}\theta_{j}(t)+\omega^{c(j)}dt+\xi_{j}\\ \mbox{(no confliction)}\\ {}\hfil\\ \Phi^{c(j),c(l)}(\theta_{j}(t),\theta_{l}(t))+\omega^{c(j)}dt+\xi_{j}\\ \mbox{(when $j$th and $l$th particles conflict)}\end{array}\right$. (97)
where $\omega^{c(j)}=\pm\omega_{0}$ represents the circular motion of chiral active matter and $c(j)=\pm 1$ represents the direction of rotation. $\Phi^{i,k}(\theta_{j}(t),\theta_{l}(t))$ determines the direction of an active particle after the collision, which depends on the chiralities of two particles; in the case of the same chirality $c(j)=c(l)$ (different chiralities $c(j)\neq c(l)$), it represents the polar (anti-polar) interaction. The model also includes the chirality flip by flipping the sign of $c(j)$ at the rate of $\gamma$.
To combine the particle-based model with the Hamiltonian of exceptional edge modes, the previous paper has further derived the hydrodynamics of chiral active matter by using the Boltzmann equations similar to Eq. (8) in Sec. 2.3.2. Compared to the Boltzmann equation of achiral active matter in Eq. (8), that of the present model has a linear chirality term $\omega^{i}\partial_{\theta}f^{i}$ with $i=\pm 1$ representing the chirality and $f^{i}$ being the distribution function of chiral active matter with the chirality $i$. The chirality flip also introduces a new linear term $\gamma(f^{-i}-f^{i})$, and the other differences are included in the collision integral (cf. Eq. (LABEL:collision_integral)). Remaining the leading order of these newly emerging terms, one obtains linear terms in its hydrodynamics that are unique to the present model. Unfortunately, the odd viscosity does not appear as a leading-order term, while it ubiquitously appears in chiral active matter (cf. Sec. 2.3.3) and plays an essential role to define the topology of active fluid as discussed in the previous section. Thus, adding an odd-viscosity term by hand and linearizing the hydrodynamic equations, one finally obtains the following effective Hamiltonian:
@@
$C=\left(\begin{array}[]{ccc}i\gamma&0&0\\ 0&i\beta&0\\ 0&0&i\beta\end{array}\right),$ (101)
where $\nu_{o}$ and $\beta$ represent the strength of the odd viscosity and the anti-poplar interaction between different chirality, respectively. The alignment interaction and chirality flips lead to the non-Hermitian coupling terms like dissipative terms discussed in Sec. 4.4. Consequently, the topological gapless modes form a pair of exceptional points as shown in Fig. 23(b). Since chirality is broadly relevant to biological active matter, such as bacteria [62] and motility assays [93] (cf. Secs. 2.1.2 and 2.1.3), exceptional edge modes should appear in such chiral active matter if one can realize the coexistence of left- and right-moving particles and their interactions.
Exceptional points can also affect the asymptotic behavior of the (dynamical) phase transition of active matter. If one considers chasers and preys on a circle, the eternally chasing (i.e., oscillatory) phase and the caught (i.e., static) phase can emerge depending on the parameters. A similar active system with nonreciprocal interactions can be realized in a collection of robots [99] (cf. Sec. 2.1.4). By analyzing such robotic active matter, the transition to the oscillatory phase accompanies exceptional points in the parameter space. The difference from the conventional phase transition emerges in, e.g., their upper critical dimension $d_{c}=8$ [302], which is different from passive systems such as the Ising model ($d_{c}=4$). Similar transitions and appearance of the exceptional points are also discussed in active Cahn-Hilliard models [303, 304, 305], which is originally a continuum model of the phase separation and has an additional nonreciprocal term that expresses activity.
@@ 5.4 Other activity-induced topological phenomena
Non-Hermiticity in active systems can further enrich topological phenomena associated with localized edge modes. A study on biological network systems [107] reveals that nonreciprocal hoppings reproduce a chiral edge current in a steady state of a stochastic process. Specifically, it considers the reaction network in Fig. 2 and numerically shows that the steady state supports the unidirectional edge current. We note that in a stochastic process with a Hermitian rate matrix (in other words, a purely dissipative stochastic process), a steady-state current is prohibited, while the non-Hermiticity due to the nonreciprocal hoppings induces the edge current. The emergence of the edge current can be schematically understood; once the particle arrives at the edge, the hopping to the edge site is much larger than that to the bulk site. Therefore, many particles stay around the edge and are transported along the direction of the hopping, which induces the chiral edge current.
@@ 6.1 Topological active matter in one and three dimensions
While topological active matter has been well explored in two-dimensional systems, topological phenomena should be also abundant in the other dimensions as we have discussed in Secs. 3 and 4. One possibility is a one-dimensional case, which can be realized by utilizing, e.g., molecular motors [30], microswimmers in a channel [308], and cell adhesions [148, 294]. In particular, while the skin effect is realized by using microchannels (cf. Sec. 5.3.1), the associated breakdown of the bulk-boundary correspondence (cf. Sec. 4.3.3) has so far not been confirmed. Due to the nonreciprocity inherent to active matter, one-dimensional active systems should offer an ideal platform to study non-Hermitian topology.
@@ 6.2 Nonlinear topology
To derive the effective Hamiltonian of active matter, we have linearized the hydrodynamic equations of active matter for the sake of simplicity (cf. Eqs. (85,86)). While the effective Hamiltonian can describe the dynamics of small fluctuations of the density and velocity fields around the steady state, the whole dynamics is ultimately governed by the nonlinear hydrodynamic equations. Recently, studies on topology of nonlinear systems have been initiated given the fact that nonlinearity is fairly common in classical systems, including photonics [310] and mechanical lattices [311]. So far, studies of nonlinear topology have revealed interplay between topology and well-known nonlinear phenomena such as topological edge solitons [312] and topological synchronization [313, 314].
@@
[…]
with $u_{j}$ being the $j$th component of the nonlinear eigenvector, $\mathbf{u}||$ being the norm of the nonlinear eigenvector, $\epsilon_{ab}$ satisfying $\epsilon_{xy}=1$, $\epsilon_{yx}=-1$, and $\epsilon_{xx}=\epsilon_{yy}=0$, and $\partial_{a,b}$ being the wavenumber derivative in the $a$, $b$ direction.
The crucial difference from linear systems is that a multiple of a nonlinear eigenvector is not necessarily a nonlinear eigenvector because of the absence of the superposition principle. Said differently, a nonlinear eigenvector also depends on its norm $\mathbf{u}||$, and one should impose $\mathbf{u}||=w$ independently of the wavenumber $\mathbf{k}$ by introducing an additional parameter $w$ of the nonlinear eigenequation. The norm dependence of the nonlinear eigenvectors and associated nonlinear topological invariants reveals the existence of the nonlinearity-induced topological phase transition, where the existence or absence of topological edge modes is tuned by the amplitude of nonlinear waves [318]. Such amplitude dependence of topological edge modes is a genuinely nonlinear effect on topological insulators and can lead to autonomous control of topological edge modes in nonlinear systems.
@@ 6.3 Topology induced by active forces
Active elements in metamaterials can induce nonreciprocal forces that are unique to nonequilibrium systems. In particular, if the metamaterial breaks the time-reversal symmetry, one can observe an active force perpendicular to the direction of the tension, which is known as odd elasticity [269]. Another time-reversal odd response is seen in odd viscosity [8], which we have already discussed in topological chiral active matter (cf. Secs. 2.3.3 and 5.2.1).
Recent studies [300, 301] have proposed topological edge modes under the non-Hermitian skin effect using odd elasticity. Since the breakdown of the time-reversal symmetry is ubiquitous in chiral active matter, one may utilize the active force to realize topological active matter. In fact, a recent experimental study [319] has observed signatures of odd elasticity and chiral edge current in collections of starfish embryos, while the topological origin of the chiral edge current is not clarified. Furthermore, since odd elasticity realizes symmetry-breaking and nonreciprocal interactions that are difficult to implement in conventional materials, such active forces can be useful for investigating non-Hermitian topological phenomena. It is also noteworthy that time-reversal odd responses are analogous to the Hall conductivity in condensed matter, and thus their direct interplay with nontrivial topology (e.g., quantization) should also be investigated.